### v1

# Personalized Heart Disease Detection via ECG Digital Twin Generation

Abstract
Heart diseases rank among the leading causes of global mortality, demonstrating a crucial need for early diagnosis and intervention. Most traditional electrocardiogram (ECG) based automated diagnosis methods are trained at population level, neglecting the customization of personalized ECGs to enhance individual healthcare management. A potential solution to address this limitation is to employ digital twins to simulate symptoms of diseases in real patients. In this paper, we present an innovative prospective learning approach for personalized heart disease detection, which generates digital twins of healthy individuals’ anomalous ECGs and enhances the model sensitivity to the personalized symptoms. In our approach, a vector quantized feature separator is proposed to locate and isolate the disease symptom and normal segments in ECG signals with ECG report guidance. Thus, the ECG digital twins can simulate specific heart diseases used to train a personalized heart disease detection model. Experiments demonstrate that our approach not only excels in generating high-fidelity ECG signals but also improves personalized heart disease detection. Moreover, our approach ensures robust privacy protection, safeguarding patient data in model development. The code can be found at https://github.com/huyjj/LAVQ-Editor
1 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_introduction
2 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_related
3 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_method_1
4 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_exp
5 Conclusion
In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals. For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor. The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information. This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development. To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level. The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation. s_conclution
Acknowledgments
This research was partially supported by National Natural Science Foundation of China under grants No. 62176231, No. 62106218, No. 82202984, No. 92259202 and No. 62132017, Zhejiang Key R&D Program of China under grant No. 2023C03053.
Appendix A: Ablation Study of threshold $l$
Figure 1: Model performance with different values of the threshold, $l$. DTW: Dynamic Time Warping; ATE: Average Treatment Effect; ECG Wise ACC: ECG Wise Accuracy.
The Vector Quantized Feature Separator incorporates textual features to navigate the dissection of features within ECG signals via a cross-attention mechanism. Within VQ-Separator, we establish a simple criterion for identifying disease symptoms in feature segments, hinging on a threshold $l$ for the cross-attention elements. Segments with values above this threshold $l$ are flagged as disease-indicative segments, while those falling below are classified as normal segments.
To assess the influence of the threshold $l$ on the authenticity of the synthesized ECGs, we conduct experiments across a spectrum of threshold values. As depicted in Figure 1, the Dynamic Time Warping (DTW)—scaled by $10^{4}$ for display purposes—remains relatively stable under varying thresholds, maintaining values below $0.6\times 10^{4}$. The ECG wise accuracy shows minimal fluctuation, consistently averaging around 0.7. In contrast, the Average Treatment Effect (ATE), represented in Figure 1 after being scaled by $10^{-1}$, exhibits more pronounced variability. Nonetheless, it sustains a level above $0.3\times 10^{-1}$, indicating a robust consistency in the validity of the results.
Appendix B: Extra Visualization Results
Referring to Figures  2,  4, and  4, we display and intuitively inspect the generated ECGs with heart diseases: Hypertrophy, Myocardial Infarction, and Conduction Disturbance, by various models. Our findings reveal:
(i) ECGs generated by StyleGAN are regularly rhythmic but are plagued by excessive noise, diminishing their authenticity. As a result, distinguishing genuine abnormalities from noise becomes a challenge, reducing their clinical utility.
(ii) PGANs, designed for personalized generation, still struggles with accurate wave depiction, hindering the identification of disease symptoms.
(iii) Our method successfully generates correct heartbeat features and markedly reduces noise, boosting ECG realism. Furthermore, our approach distinctly reveals clear pathological signs. As indicated by orange rectangles in Figure  2, we observe T-wave flattening in lead I, indicative of potential hypertrophic cardiomyopathy. Figure  4 display pronounced QRS abnormalities in lead V3, where the T wave is notably higher than the R and S waves. Figure  4 shows instances of the P wave failing to progress to the ventricle, potentially indicative of ventricular or junctional escape rhythms.
Figure 2: Visualization of generated ECGs with the Hypertrophy (HYP) heart disease patterns by different models. For better viewing, we only display three leads of ECGs: I, aVR, and V3.
Figure 3: Visualization of generated ECGs with the Myocardial Infarction (MI) heart disease patterns by different models. For better viewing, we only display three leads of ECGs: I, aVR, and V3.
Figure 4: Visualization of generated ECGs with the Conduction Disturbance (CD) heart disease patterns by different models. For better viewing, we only display three leads of ECGs: I, aVR, and V3.

### v2

@@
 Moreover, our approach ensures robust privacy protection, safeguarding patient data in model development.
-The code can be found at https://github.com/huyjj/LAVQ-Editor
-1 Conclusion
-In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals.
-For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor.
-The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information.
-This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development.
-To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level.
-The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation.
-s_introduction
-2 Conclusion
-In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals.
-For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor.
-The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information.
-This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development.
-To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level.
-The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation.
-s_related
-3 Conclusion
-In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals.
-For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor.
-The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information.
-This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development.
-To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level.
-The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation.
-s_method_1
-4 Conclusion
-In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals.
-For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor.
-The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information.
-This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development.
-To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level.
-The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation.
-s_exp
-5 Conclusion
-In this paper, we introduce a novel prospective learning framework for creating personalized ECG digital twins that simulate the heart conditions of diagnosed individuals.
-For the purpose of realistic ECG generation, we propose a location-aware ECG digital twin generation model based on vector quantization, termed LAVQ-Editor.
-The model skillfully segregates and edits normal and disease-indicative features in ECGs, preserving individual characteristics while injecting target disease information.
-This enhances understanding of personalized heart diseases prospectively while safeguarding privacy in model development.
-To the best of our knowledge, our LAVQ-Editor is the first framework for personalized ECG digital twin generation in contrast to previous work that only generated data at population level.
-The versatility of our approach extends beyond diagnostics, holding potential for application in ethically-conducted scientific research, such as clinical trial simulation.
-s_conclution
+The code can be found at https://github.com/huyjj/LAVQ-Editor.
 Acknowledgments
-This research was partially supported by National Natural Science Foundation of China under grants No.
-62176231, No.
-62106218, No.
-82202984, No.
-92259202 and No.
-62132017, Zhejiang Key R&D Program of China under grant No.
-2023C03053.
-Appendix A: Ablation Study of threshold $l$
+This research was supported by Zhejiang Provincial Natural Science Foundation of China (LY23F020019), National Natural Science Foundation of China (62106218 & 82200558), Excellent Young Scientists Fund (82122036), Basic and Applied Basic Research Foundation of Guangdong Province (2022A1515110722), and Guangdong Provincial Medical Science & Technology Research Fund Project (A2023011).
+Appendix A: Pseudocode of Training
+  Initialize VQ-Separator, Generator $G$, and Discriminator $D$
+  for each training epoch do
+     for each batch in the dataset do
+        $X_{ecg}\leftarrow$ pre-diagnosis ECG, $X_{text}\leftarrow$ the report of ECG
+        $X_{ecg}^{t}\leftarrow$ reference ECG, $X_{text}^{t}\leftarrow$ the report of reference ECG
+        Optimizer Discriminator $D$:
+        $f_{d},f_{p}\leftarrow\text{VQ-Separator}({X}_{ecg},X_{text})$
+        $f_{d}^{t},f_{p}^{t}\leftarrow\text{VQ-Separator}({X}_{ecg}^{t},X_{text}^{t})$
+        $\hat{X_{ecg}^{t}}\leftarrow G_{dec}(f_{p},f_{d}^{t})$
+        $\hat{X}_{ecg}^{rec}\leftarrow G_{dec}(f_{p},f_{d})$
+        $f_{d}^{G},f_{p}^{G}\leftarrow\text{VQ-Separator}(\hat{X}_{ecg}^{t},X_{text}^{t})$
+        $\hat{f_{d}},D_{real}\leftarrow D(X_{ecg})$
+        $\hat{f_{d}^{rec}},D_{fake}\leftarrow D(\hat{X}_{ecg}^{rec})$
+        Calculate losses:
+        $L_{adv}^{D}\leftarrow\log(D_{real})+\log(1-D_{fake})$
+        $L_{sim}^{D}\leftarrow\mathcal{L}_{sim}^{D}=1-\frac{f_{d}^{D}\cdot f_{d}}{\parallel f_{d}^{D}\parallel\parallel f_{d}\parallel}$
+        Update $D$ by minimizing $L_{sim}^{D}+L_{adv}^{D}$
+        Optimizer Generator $G$:
+        $f_{d},f_{p},L_{vq}\leftarrow\text{VQ-Separator}({X}_{ecg},X_{text})$
+        $f_{d}^{t},f_{p}^{t}\leftarrow\text{VQ-Separator}({X}_{ecg}^{t},X_{text}^{t})$
+        $\hat{X_{ecg}^{t}}\leftarrow G_{dec}(f_{p},f_{d}^{t})$
+        $\hat{X_{ecg}^{rec}}\leftarrow G_{dec}(f_{p},f_{d})$
+        $f_{d}^{G},f_{p}^{G}\leftarrow\text{VQ-Separator}(\hat{X}_{ecg}^{t},X_{text}^{t})$
+        $\hat{f_{d}^{rec}},D_{fake}\leftarrow D(\hat{X}_{ecg}^{rec})$
+        Calculate losses:
+        $L_{recon}\leftarrow||X_{ecg}-\hat{X}_{ecg}^{rec}||^{2}$
+        $L_{adv}\leftarrow\log(1-D_{fake})$
+        Update $G$ by minimizing $L_{recon}+L_{vq}+L_{adv}^{G}+L_{sim}^{G}$
+     end for
+     Evaluate model performance on validation set
+  end for
+Algorithm 1 Training Process of the LAVQ-Editor
+Appendix B: Ablation Study of threshold $l$
 Figure 1: Model performance with different values of the threshold, $l$. DTW: Dynamic Time Warping; ATE: Average Treatment Effect; ECG Wise ACC: ECG Wise Accuracy.
@@
 Nonetheless, it sustains a level above $0.3\times 10^{-1}$, indicating a robust consistency in the validity of the results.
-Appendix B: Extra Visualization Results
-Referring to Figures  2,  4, and  4, we display and intuitively inspect the generated ECGs with heart diseases: Hypertrophy, Myocardial Infarction, and Conduction Disturbance, by various models.
+Appendix C: Extra Visualization Results
+Referring to Figures  3,  5, and  5, we display and intuitively inspect the generated ECGs with heart diseases: Hypertrophy, Myocardial Infarction, and Conduction Disturbance, by various models.
 Our findings reveal:
@@
 Furthermore, our approach distinctly reveals clear pathological signs.
-As indicated by orange rectangles in Figure  2, we observe T-wave flattening in lead I, indicative of potential hypertrophic cardiomyopathy.
-Figure  4 display pronounced QRS abnormalities in lead V3, where the T wave is notably higher than the R and S waves.
-Figure  4 shows instances of the P wave failing to progress to the ventricle, potentially indicative of ventricular or junctional escape rhythms.
+As indicated by orange rectangles in Figure  3, we observe T-wave flattening in lead I, indicative of potential hypertrophic cardiomyopathy.
+Figure  5 display pronounced QRS abnormalities in lead V3, where the T wave is notably higher than the R and S waves.
+Figure  5 shows instances of the P wave failing to progress to the ventricle, potentially indicative of ventricular or junctional escape rhythms.
+For timing problems, as shown in Figure 2, our model can scale the signals based on the statistical RR interval length after morphological editing.
+Figure 2: Example of ECG described as “sinus tachycardia apart from rate, normal ecg.” generated from the real normal ECG.
 Figure 3: Visualization of generated ECGs with the Hypertrophy (HYP) heart disease patterns by different models. For better viewing, we only display three leads of ECGs: I, aVR, and V3.